# Deconstructing The Ethics of Large Language Models from Long-standing Issues to New-emerging Dilemmas

### v1

@@ Abstract
Large Language Models (LLMs) have achieved unparalleled success across diverse language modeling tasks in recent years. However, this progress has also intensified ethical concerns, impacting the deployment of LLMs in everyday contexts. This paper provides a comprehensive survey of ethical challenges associated with LLMs, from longstanding issues such as copyright infringement, systematic bias, data privacy, to emerging problems like truthfulness and social norms. We critically analyze existing research aimed at understanding, examining, and mitigating these ethical risks. Our survey underscores integrating ethical standards and societal values into the development of LLMs, thereby guiding the development of responsible and ethically aligned language models.
@@ 1 Introduction
In the past few years, the field of artificial intelligence (AI) has witnessed a surge in the development of large language models (LLMs). These advanced computational language models have demonstrated remarkable performance across a spectrum of language modeling tasks [46, 254, 287, 338, 348, 347, 188]. Their capabilities are exemplified in natural language generation [38, 47, 205], where they can create coherent and contextually relevant text, question answering [15, 331, 351], where they effectively retrieve or infer information in response to queries, and complex reasoning tasks [118, 130, 328, 305], which involve navigating through intricate problem-solving processes. Despite these advancements, LLMs have also raised substantial ethical concerns. As these models become increasingly integrated into daily life, addressing these ethical challenges becomes paramount. The concerns are multifaceted, encompassing issues such as privacy [302], copyright, robustness [329], bias, and the potential for misuse. Given their ability to understand and generate human-like responses, there’s a growing discourse on ensuring these responses are not only accurate but also ethically aligned with societal norms and values.
@@ 2.1.1 Privacy: Issues and Challenges
The debut of ChatGPT marked a pivotal moment, sparking a surge in AI research, an increase in new ventures, and widespread adoption of Large Language Models (LLMs). However, there is an emerging agreement that, while LLMs offer substantial advantages, they also present notable societal challenges, especially privacy. In this section, we first introduce issues and potential challenges and then discuss major solutions regarding these issues (e.g. Section 2.1.2 deferentially private LLMs and other emerging techniques in Section 2.1.3 The concerns in privacy could be mainly summarized in twofold, memorization and privacy attacks.
Memorization.
@@
All machine-learning (ML) models, including LLMs, inherently memorize to some extent, as they learn by observing and recalling training data. However, this problem becomes severe when it comes to LLMs because of its tremendous size and capacity. We list the main aspects of risk factors that may affect the memorization issue.
@@
  • Model size: The capacity of a model significantly impacts its memorization ability. Larger models, as shown by [44] and [263], tend to memorize more data and do so at a faster rate. This memorization is not directly linked to model performance, as shown by comparing GPT-2 and GPT-Neo models. The trend suggests that neural networks’ capacity to memorize is substantial and growing, outpacing the size increase of language datasets.
@@
  • Size of the dataset: Research on dataset size and memorization reveals contrasting findings. Li et al. discovered that larger datasets lead to less memorization, evidenced by a decline in canary extraction success over training time [154]. Conversely, Biderman et al. found that points memorized early in training tend to be retained in fully trained models, suggesting persistent memorization despite dataset size [30].
@@
  • Data duplication is a key factor in memorization for Large Language Models (LLMs). Lee et al. [149] observed that data duplication in large web datasets follows a power law, with a small fraction of data being highly duplicated. This duplication significantly increases memorization, as models trained on deduplicated datasets exhibit much lower rates of outputting memorized text. Kandpal et al. [137] further demonstrated that sequences repeated in the dataset are generated far more frequently by LLMs. Despite this, memorization still occurs even with little or no data duplication, indicating other contributing factors to memorization beyond mere duplication.
@@
  • Prompt length and type significantly affect memorization in Large Language Models (LLMs). Mccoy et al. observed that longer generated sequences (n) tend to produce more novel content, reducing memorization [192]. Conversely, longer prompts increase memorization for a constant n, as shown by [44]. Additionally, specific token types, like nouns and numbers, are memorized faster than others, such as verbs and adjectives. Kharitonov et al. found that larger subword vocabularies in tokenizers lead to increased memorization, possibly due to reduced sequence length making it easier for models to memorize [140].
Privacy Attack.
@@
The robustness of Large Language Models (LLMs) may be weakened by privacy attacks. We list three scenarios that may bring privacy risks to the robustness issues of LLMs as follows.
@@
  • Training data extraction from language models (LMs) is a privacy attack enabling adversaries to retrieve sensitive data using query access. Carlini et al. pioneered this method, involving generating candidate targets, applying a membership inference attack (MIA), and selecting top-k candidates [43]. Their experiments on GPT-2 demonstrated the feasibility of extracting training data, including sensitive personal information. Subsequent research by [320, 335] introduced improvements in candidate generation and MIA processes, significantly enhancing extraction precision. Nasr et al. extended these attacks to production LMs like ChatGPT and open-source models, revealing higher memorization levels than previously understood [210]. This line of research underscores the potential privacy risks inherent in LMs and the effectiveness of training data extraction attacks.
@@ 2.1.2 Differentially Private LLMs
Differential privacy (DP) [71] emerges as the primary scheme to address data privacy concerns. Acknowledged as de facto golden standard, differential privacy provides mathematical rigor to the algorithms involving sensitive information to be protected. Essentially, an algorithm is differentially private if the output distribution is relatively close, tailored by certain privacy parameters whether an individual’s data is present or not in the dataset. More formally,
@@
Pre-LLM Era.
We first explore existing methods that employ DP training, where a language model is usually trained from scratch using variants of DP-SGD. An early attempt, DP-FedAve [193] dates back to the ante-transformer era. It targets recurrent language models and introduces a DP optimization technique inspired by a federated averaging algorithm. Consequently, differential privacy is defined on the user level. To improve the privacy-utility trade-off, a later work, Selective DP-SGD [246] introduces the concept of selective differential privacy, which provides focused protection for sensitive attributes only in one training example. Note that this method only applies to RNN-based language models. Moving forward to pre-trained transformer language models, two closely related works [111, 10] improve DP-SGD and train BERT with DP guarantees. Both consider the protection level as item-level, which is one training example containing several words. The latter work [10] focuses on training heuristics that bring more efficiency and can be implemented on BERT-large.
@@
LLM Era.
LLMs demonstrate compelling capabilities such as in-context learning merely within the inference stage. Privacy-preserving approaches lying in this category bypass the projection of DP-SGD and commonly add perturbation to more accessible information sources such as prompts or embeddings, leaving LLMs parameters frozen. With respect to in-context learning, two works [299, 260] emerge with a similar scheme of ‘divide-and-privately-aggregate’, however, considering different privacy levels. DP-ICL [299] aggregates the LLM responses for each group of exemplars with differential privacy. Two mechanisms are proposed for private aggregation: embedding space aggregation and keyword space aggregation. DP-ICL is on the user level while the later work [260] zooms into the example level, the aggregation algorithms are based on the Gaussian mechanism and exponential mechanism and applied to exemplars in sensitive datasets. Another work on privacy-preserving prompt tuning called RAPT [157] also privatizes source datasets with DP guarantees, where tokens are reconstructed with randomized mechanisms, and then trained jointly with the downstream tasks. Last, we include three recent methods that apply DP by adding perturbation to embeddings. DP-forward [69] devises an analytic matrix Gaussian mechanism that perturbs the embedding matrices in the forward pass of language models. Split-N-Denoise [185] further provides a framework where the embeddings are first perturbed on the user side and then transmitted to the server. A denoising module can be trained to produce outputs given noisy responses from the server LLMs. Both works consider local DP. Shortly after, InferDPT [265] moves to document-level DP that protects sensitive information in prompts for black-box LLM inference. The proposed pipeline contains a perturbation module based on an exponential mechanism and an extraction module that selects coherent and consistent text from the perturbed generation result.
@@ 2.2 Copyright
Copyright has been a long-existing legal issue in the natural language domain [23] that calls for research on encoding imperceptible and indelible signatures on plain texts to protect the property of authorship [8]. In literature, as an information hiding application [22], the traditional techniques extend from steganography [59] to watermarking [248]. In the language model era, copyrights preserving techniques further develop to protect the model rather than sorely the data, where backdoor [50, 88, 158, 80, 184] and watermark [143] are two main streams.
@@ 2.2.1 Backdoor
Backdoor attacks [88, 212, 136, 78, 16, 183, 301, 182] inject poisoned samples containing a specific trigger into the training dataset. Models trained on it will predict the attacker-specific target label when they encounter samples with the trigger during inference while behaving normally when the trigger is absent. By embedding a unique trigger pattern within a model through a backdoor, a unique relationship between the trigger and the target label is established. The presence of samples with the trigger will consistently induce the model to predict the corresponding target label, which can be used to signify the model’s ownership or origin, particularly in situations where the model is not accessible, such as in a black-box setting.
Pre-LLM Era.
@@
Adi et al. first introduce that Backdoor can be used as watermarks for ownership verification [6]. To avoid detection, Xiang et al. propose a semantic and robust watermarking scheme for natural language generation (NLG) models that utilize unharmful phrase pairs as watermarks for intellectual property (IP) protection [300]. He et al. use lexical replacements of specific words to demonstrate ownership for LLMs deployed through APIs [106]. In addition, large pre-trained language models (PLMs) require fine-tuning on downstream datasets, which makes it hard to claim the ownership of PLMs. Gu et al. show that PLMs can be watermarked with a multi-task learning framework by embedding backdoors, making watermarks difficult to remove even after fine-tuning the models on multiple tasks [87]. Liu et al. present a novel watermarking technique using a backdoor-based membership inference approach via marking a small subset of samples for data copyright protection in the black-box setting [178].
LLM Era.
@@
Copyright protection of LLMs has become crucial due to the substantial training cost associated with these models. EmbMarker [224] proposes to implant backdoors on embeddings of LLMs. Specifically, it selects moderately frequent words as triggers, defines a target embedding as the watermark, and uses a backdoor function to embed it. Lucas et al. propose an attack to identify trigger words or phrases by analyzing open-ended generations from LLMs with backdoor watermarks [181]. It is shown that triggers based on random common words are easier to detect than those based on rare tokens.
Discussion.
@@
We suggest that the exploration of stricter settings is necessary. For example, in most research, data owners have access to the percentage of their data within the total training set, which necessitates knowledge of tasks associated with PLMs. Hence, how to adapt the backdoor-based methods for stricter settings in copyright protection remains an open direction. In addition, as the field of backdoor-based copyright protection advances, an increasing number of tailored model-stealing techniques are being studied, such as knowledge distillation [109]. It is essential to explore the resilience of backdoor-based algorithms against potential attacks that adversaries may employ. Finally, the effectiveness of backdoor-based copyright protection for LMs still lacks a comprehensive theoretical framework. The clarity of such a framework remains an open question in this field.
@@ 2.2.2 Watermark
Watermark is first concerned as an information hiding technology for a small amount of information [227]. Mir et al. apply this technique to protect the copyright of web content [201]. Early approaches of watermarking include text-meaning representations of sentences for information hiding by syntactic rules [12], watermarking on the format of documents by vertical and horizontal line-shifting [37], watermarking by inserting zero-width control characters in Hyper Text Markup Language (HTML) [9], watermarking on semantics by synonyms substitution [24, 266], and zero-watermarks by using word length [126] and contents of text [127].
LLM Era.
@@
Watermarking at the current stage focuses more on the model schemes for watermarked generation. As pioneers, Kirchenbauer et al. propose an LLM watermarking algorithm by adding token-level bias in the decoding stage [143]. Kuditipudi et al. design a distortion-free watermark to preserve the original distribution of LLM during watermarking [146]. Ren et al. consider the semantic embedding in hashing tokens [241] and Fu et al. concern semantic word similarity to enhance the robustness [75]. Yoo et al. embed multi-bit information into the watermark, which succeeds traditional steganography [316]. They inject the watermark via word replacement after initial generation, which is further integrated into one stage by [283]. Christ et al. propose a computationally undetectable watermark theoretically if the secret key is inaccessible [54]. Liu et al. propose a private watermark utilizing separated neural networks respectively for generation and detection [165]. The aforementioned works focus more on the token level, while there are emerging works focusing on a higher-level perspective. Hou et al. introduce a sentence-level semantic watermark that aims at periphrastic robustness [112, 113]. For applications, some works mention the importance of watermarking the ownership of datasets via inference [186, 170]. Yao et al. introduce copyright protection for prompts via watermarking [309].
Discussion.
@@
One of the main challenges for watermarking is its popularization and the opening of corresponding detection methods and configurations. Hopefully, this requires administrative oversight from government and industry associations. US Federal, China, and Europe have mentioned potential proposals in some of the government documents, e.g., Interim Measures for Generative Artificial Intelligence Service Management of China, Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence of the US, and the European Union’s AI Act. Moreover, the definition and notion of authorship are slightly ambiguous as human-LLM collaboration and multi-agent generation are becoming mainstream. Tripto et al. discover that literate studies have contrasting perspectives on whether authorship remains the same after paraphrasing, as paraphrasing deviates the style of text dramatically [271]. Meanwhile, further improvement on the watermark’s robustness to attack [285], generalization to short contents, reduction of impact on text quality, and differentiation to direct machine-generated text detection [82, 203, 175, 173] are worth exploring.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved unparalleled success across diverse language modeling tasks in recent years. However, this progress has also intensified ethical concerns, impacting the deployment of LLMs in everyday contexts. This paper provides a comprehensive survey of ethical challenges associated with LLMs, from longstanding issues such as copyright infringement, systematic bias, and data privacy, to emerging problems like truthfulness and social norms. We critically analyze existing research aimed at understanding, examining, and mitigating these ethical risks. Our survey underscores integrating ethical standards and societal values into the development of LLMs, thereby guiding the development of responsible and ethically aligned language models.
keywords:
Large Language Models, Trustworthy, Ethics, Survey
@@ 1 Introduction
In the past few years, the field of artificial intelligence (AI) has witnessed a surge in the development of large language models (LLMs). These advanced computational language models have demonstrated remarkable performance across a spectrum of language modeling tasks [46, 259, 292, 344, 354, 353, 192, 312]. Their capabilities are exemplified in natural language generation [38, 47, 209], where they can create coherent and contextually relevant text, question answering [15, 337, 357], where they effectively retrieve or infer information in response to queries, and complex reasoning tasks [119, 131, 334, 310], which involve navigating through intricate problem-solving processes. Despite these advancements, LLMs have also raised substantial ethical concerns. As these models become increasingly integrated into daily life, addressing these ethical challenges becomes paramount. The concerns are multifaceted, encompassing issues such as privacy [307], copyright, robustness [335], bias, and the potential for misuse. Given their ability to understand and generate human-like responses, there’s a growing discourse on ensuring these responses are not only accurate but also ethically aligned with societal norms and values.
@@ 2.1.1 Privacy: Issues and Challenges
Data privacy has long been a concern, but there is a growing consensus that while Large Language Models (LLMs) offer impressive capabilities, they also raise significant data privacy issues today. In this section, we first introduce issues and potential challenges and then discuss major solutions regarding these issues (e.g. Section 2.1.2 deferentially private LLMs and other emerging techniques in Section 2.1.3). The concerns in privacy could be mainly summarized in twofold, memorization and privacy attacks as illustrated in Figure 3.
Figure 3: Data privacy issues & challenges detailed categories and mitigation methods.
@@
Memorization. All machine-learning (ML) models, including LLMs, inherently memorize to some extent, as they learn by observing and recalling training data. However, this problem becomes severe when it comes to LLMs because of its tremendous size and capacity. We list the main aspects of risk factors that may affect the memorization issues.
@@
  • Model capacity: The capacity of a model significantly impacts its memorization ability. Larger models, as shown by [44] and [268], tend to memorize more data and do so at a faster rate. This memorization is not directly linked to model performance, as shown by comparing GPT-2 and GPT-Neo models. The trend suggests that neural networks’ capacity to memorize is substantial and growing, outpacing the size increase of language datasets.
@@
  • Dataset scale: Research on dataset size and memorization reveals contrasting findings. Li et al. [156] discovered that larger datasets lead to less memorization, evidenced by a decline in canary extraction success over training time. Conversely, Biderman et al. [30] found that points memorized early in training tend to be retained in fully trained models, suggesting persistent memorization despite dataset size.
@@
  • Sample duplication is a key factor in memorization for Large Language Models (LLMs). Lee et al. [150] observed that data duplication in large web datasets follows a power law, with a small fraction of data being highly duplicated. This duplication significantly increases memorization, as models trained on deduplicated datasets exhibit much lower rates of outputting memorized text. Kandpal et al. [138] further demonstrated that sequences repeated in the dataset are generated far more frequently by LLMs. Despite this, memorization still occurs even with little or no data duplication, indicating other contributing factors to memorization beyond mere duplication.
@@
  • Prompting significantly affects memorization in Large Language Models (LLMs). Mccoy et al. [196] observed that longer generated sequences tend to produce more novel content, reducing memorization. Conversely, longer prompts increase memorization for a constant n, as shown by [44]. Additionally, specific token types, like nouns and numbers, are memorized faster than others, such as verbs and adjectives. Kharitonov et al. found that larger subword vocabularies in tokenizers lead to increased memorization, possibly due to reduced sequence length making it easier for models to memorize [141].
@@
Privacy Attack. The robustness of Large Language Models (LLMs) may be weakened by privacy attacks. We list three scenarios that may bring privacy risks to the robustness issues of LLMs as follows.
@@
  • Training data extraction is a privacy attack enabling adversaries to retrieve sensitive data using query access. Carlini et al. pioneered this method, involving generating candidate targets, applying a membership inference attack (MIA), and selecting top-k candidates [43]. Their experiments on GPT-2 demonstrated the feasibility of extracting training data, including sensitive personal information. Subsequent research by [326, 341] introduced improvements in candidate generation and MIA processes, significantly enhancing extraction precision. Nasr et al. extended these attacks to production LMs like ChatGPT and open-source models, revealing higher memorization levels than previously understood [214]. This line of research underscores the potential privacy risks inherent in LMs and the effectiveness of training data extraction attacks.
@@ 2.1.2 Differentially Private LLMs
Differential privacy (DP) [72] emerges as the primary scheme to address data privacy concerns. Acknowledged as de facto golden standard, differential privacy provides mathematical rigor to the algorithms involving sensitive information to be protected. Essentially, an algorithm is differentially private if the output distribution is relatively close, tailored by certain privacy parameters whether an individual’s data is present or not in the dataset. More formally, we denote differential privacy as follows.
@@
Pre-LLM Era. We first explore existing methods that employ DP training, where a language model is usually trained from scratch using variants of DP-SGD. An early attempt, DP-FedAve [197] dates back to the ante-transformer era. It targets recurrent language models and introduces a DP optimization technique inspired by a federated averaging algorithm. Consequently, differential privacy is defined on the user level. To improve the privacy-utility trade-off, a later work, Selective DP-SGD [250] introduces the concept of selective differential privacy, which provides focused protection for sensitive attributes only in one training example. Note that this method only applies to RNN-based language models. Moving forward to pre-trained transformer language models, two closely related works [112, 10] improve DP-SGD and train BERT with DP guarantees. Both consider the protection level as item-level, which is one training example containing several words. The latter work [10] focuses on training heuristics that bring more efficiency and can be implemented on BERT-large.
@@
LLM Era. LLMs demonstrate compelling capabilities such as in-context learning merely within the inference stage. Privacy-preserving approaches lying in this category bypass the projection of DP-SGD and commonly add perturbation to more accessible information sources such as prompts or embeddings, leaving LLMs parameters frozen. With respect to in-context learning, two works [304, 265] emerge with a similar scheme of ‘divide-and-privately-aggregate’, however, considering different privacy levels. DP-ICL [304] aggregates the LLM responses for each group of exemplars with differential privacy. Two mechanisms are proposed for private aggregation: embedding space aggregation and keyword space aggregation. DP-ICL is on the user level while the later work [265] zooms into the example level, the aggregation algorithms are based on the Gaussian mechanism and exponential mechanism and applied to exemplars in sensitive datasets. Another work on privacy-preserving prompt tuning called RAPT [159] also privatizes source datasets with DP guarantees, where tokens are reconstructed with randomized mechanisms, and then trained jointly with the downstream tasks. Last, we include three recent methods that apply DP by adding perturbation to embeddings. DP-forward [70] devises an analytic matrix Gaussian mechanism that perturbs the embedding matrices in the forward pass of language models. Split-N-Denoise [189] further provides a framework where the embeddings are first perturbed on the user side and then transmitted to the server. A denoising module can be trained to produce outputs given noisy responses from the server LLMs. Both works consider local DP. Shortly after, InferDPT [270] moves to document-level DP that protects sensitive information in prompts for black-box LLM inference. The proposed pipeline contains a perturbation module based on an exponential mechanism and an extraction module that selects coherent and consistent text from the perturbed generation result.
@@ 2.2 Copyright
Copyright has been a long-existing legal issue in the natural language domain [23] that calls for research on encoding imperceptible and indelible signatures on plain texts to protect the property of authorship [8]. In literature, as an information hiding application [22], the traditional techniques extend from steganography [59] to watermarking [253]. In the language model era, copyrights preserving techniques further develop to protect the model rather than sorely the data, where backdoor [50, 89, 62, 160, 81, 188] and watermark [144] are two main streams. The hierarchy in this section is portrayed in Figure 4.
Figure 4: Copyright methods.
@@ 2.2.1 Backdoor
In backdoor attacks [89, 216, 137, 79, 16, 186, 306, 185, 187], the attacker constructs poisoned samples by adding an attacker-defined trigger to a fraction of the training samples and changing the associated labels to a specific target class. A backdoor can be injected by training the model with a mixture of clean and poisoned samples. A backdoor-compromised model functions normally with clean inputs but exhibits abnormal behavior when presented with inputs containing a specific trigger. By embedding a unique trigger pattern within a model through a backdoor, a distinct relationship between the trigger and the target label is established. In classification tasks, the presence of the trigger will consistently induce the model to predict the corresponding target label. These properties can be used to signify the model’s ownership or origin, particularly in situations where the model is not accessible, such as in a black-box setting.
@@
Pre-LLM Era. Adi et al. first introduce that Backdoor can be used as watermarks for ownership verification [6]. To avoid detection, Xiang et al. propose a semantic and robust watermarking scheme for natural language generation (NLG) models that utilize unharmful phrase pairs as watermarks for intellectual property (IP) protection [305]. He et al. use lexical replacements of specific words to demonstrate ownership for LLMs deployed through APIs [107]. In addition, large pre-trained language models (PLMs) require fine-tuning on downstream datasets, which makes it hard to claim the ownership of PLMs. Gu et al. show that PLMs can be watermarked with a multi-task learning framework by embedding backdoors, making watermarks difficult to remove even after fine-tuning the models on multiple tasks [88]. Shokri et al. investigate membership inference attacks on machine learning models trained by commercial ”machine learning as a service” providers such as Google and Amazon, determining if a data record was part of the training dataset. [252]. Liu et al. present a novel watermarking technique using a backdoor-based membership inference approach via marking a small subset of samples for data copyright protection in the black-box setting [180].
@@
LLM Era. Copyright protection of LLMs has become crucial due to the substantial training cost associated with these models. Liu et al. indicate that LLMs are vulnerable to model extraction attacks, wherein attackers can copy the model using query texts and returned embeddings, potentially building their own LLMs and causing significant losses for the original model owners [181]. EmbMarker [228] proposes to implant backdoors on embeddings of LLMs. Specifically, it selects moderately frequent words as triggers, defines a target embedding as the watermark, and uses a backdoor function to embed it. Lucas et al. propose an attack to identify trigger words or phrases by analyzing open-ended generations from LLMs with backdoor watermarks [184]. It is shown that triggers based on random common words are easier to detect than those based on rare tokens.
@@
Discussion. We suggest that the exploration of stricter settings is necessary. For example, in most research, data owners have access to the percentage of their data within the total training set, which necessitates knowledge of tasks associated with PLMs. Hence, how to adapt the backdoor-based methods for stricter settings in copyright protection remains an open direction. In addition, as the field of backdoor-based copyright protection advances, an increasing number of tailored model-stealing techniques are being studied, such as knowledge distillation [110]. It is essential to explore the resilience of backdoor-based algorithms against potential attacks that adversaries may employ. Finally, the effectiveness of backdoor-based copyright protection for LMs still lacks a comprehensive theoretical framework. The clarity of such a framework remains an open question in this field.
@@ 2.2.2 Watermark
Pre-LLM Era. Watermark is first concerned as an information hiding technology for a small amount of information [231]. Mir et al. apply this technique to protect the copyright of web content [205]. Early approaches of watermarking include text-meaning representations of sentences for information hiding by syntactic rules [12], watermarking on the format of documents by vertical and horizontal line-shifting [37], watermarking by inserting zero-width control characters in Hyper Text Markup Language (HTML) [9], watermarking on semantics by synonyms substitution [24, 271], and zero-watermarks by using word length [127] and contents of text [128].
@@
LLM Era. Watermarking at the current stage focuses more on the model schemes for watermarked generation. As pioneers, Kirchenbauer et al. propose an LLM watermarking algorithm by adding token-level bias in the decoding stage [144]. Kuditipudi et al. design a distortion-free watermark to preserve the original distribution of LLM during watermarking [147]. Ren et al. consider the semantic embedding in hashing tokens [245] and Fu et al. concern semantic word similarity to enhance the robustness [76]. Yoo et al. embed multi-bit information into the watermark, which succeeds traditional steganography [322]. They inject the watermark via word replacement after initial generation, which is further integrated into one stage by [288]. Christ et al. propose a computationally undetectable watermark theoretically if the secret key is inaccessible [54]. Liu et al. propose a private watermark utilizing separated neural networks respectively for generation and detection [167]. The aforementioned works focus more on the token level, while there are emerging works focusing on a higher-level perspective. Hou et al. introduce a sentence-level semantic watermark that aims at periphrastic robustness [113, 114]. For applications, some works mention the importance of watermarking the ownership of datasets via inference [190, 172]. Yao et al. introduce copyright protection for prompts via watermarking [315].
@@
Discussion. One of the main challenges for watermarking is its popularization and the opening of corresponding detection methods and configurations. Hopefully, this requires administrative oversight from government and industry associations. US Federal, China, and Europe have mentioned potential proposals in some of the government documents, e.g., Interim Measures for Generative Artificial Intelligence Service Management of China, Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence of the US, and the European Union’s AI Act. Moreover, the definition and notion of authorship are slightly ambiguous as human-LLM collaboration and multi-agent generation are becoming mainstream. Tripto et al. discover that literate studies have contrasting perspectives on whether authorship remains the same after paraphrasing, as paraphrasing deviates the style of text dramatically [276]. Meanwhile, further improvement on the watermark’s robustness to attack [290], generalization to short contents, reduction of impact on text quality, and differentiation to direct machine-generated text detection [83, 207, 177, 175] are worth exploring.